OpenStudy (mackenzie2013):

Which of the following sets of numbers is a Pythagorean triple? 42, 40, 82 42, 40, 58 41, 41, 58 40, 41, 57 Wouldnt C be corrrecT?

ganeshie8 (ganeshie8):

pythagorean triples satisfy below : \(\large a^2 + b^2 = c^2\)

ganeshie8 (ganeshie8):

does C satisfy that relation ?

OpenStudy (mackenzie2013):

i got 57.9? wouldnt you round it or no?

ganeshie8 (ganeshie8):

no rounding allowed, try other options :)

OpenStudy (mackenzie2013):

So D!

ganeshie8 (ganeshie8):

hmm i don't think so, check again

ganeshie8 (ganeshie8):

Is this true \(\large 40^2 + 41^2 = 57^2\) ?

OpenStudy (mackenzie2013):

Its b

ganeshie8 (ganeshie8):

Yep !